OpenStudy (anonymous):

sin(sin^{-1} \frac{ -sqrt{3} }{ 2 }+cos^{-1} \frac{ 1 }{ 2 })

OpenStudy (ash2326):

@klbala2006 Do you know the range of \(\sin^{-1} x\) ?

OpenStudy (campbell_st):

looking at it its all 4th quadrant 

OpenStudy (anonymous):

i need to evaluate this function . @ash2326 no i dont

OpenStudy (ash2326):

Do you know the range and domain of sin (x) ?

OpenStudy (anonymous):

@ash2326 domain of sin(x) is (R), while its range is [-1,1]

OpenStudy (ash2326):

Good, to have the inverse of sin function, we limit it's range domain= range of sin x= [-1, 1] Range = \(\large [\frac{-\pi}{2}, \frac{\pi}{2}]\) DO you get this?

OpenStudy (anonymous):

yes i do

OpenStudy (ash2326):

if x is positive, then y lies in \([0, \frac {\pi}{2}]\) if x is negative, then y lies in \([\frac{-\pi}{2}, 0)\) Now what's \(\sin^{-1}(\frac{-\sqrt 3}{2})\) ?

OpenStudy (anonymous):

\[-45\]

OpenStudy (anonymous):

\[-\frac{ \pi }{ 4 }\] in radian

OpenStudy (anonymous):

@ash2326 complete please

OpenStudy (ash2326):

Sorry, I was away. But sin (-45) is not \(\frac{-\sqrt 3}{2}\) it's, \(\frac{-1}{\sqrt 2}\).